# Never-Ending Behavior-Cloning Agent for Robotic Manipulation

### v2

@@ 3.3 Skill-Shared Representation Distillation Module
where $\hat{B}=\sum_{j=1}^{B}\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}}$, and $\tau$ is a hyper-parameter represents distillation temperature.
3.4 Skill-Specific Evolving Planner
@@
where $\mathcal{B}\in\mathbb{R}^{N^{b}\times D^{s}}$ is initialized by $N^{b}$ zero vectors. $\mathcal{C}\in\mathbb{R}^{N^{b}}$ is the cosine similarity matrix between the sentence information $\mathbf{l}_{s}$ and each vector in $\mathcal{B}$. if $\mathcal{C}_{max}>\delta$, we set $h=\mathrm{argmax}(\mathcal{C})$; otherwise, we set $h=\mathrm{nonzero}(\mathcal{B})+1$ and $\mathcal{C}_{max}=1$, where $\delta$ is a fixed threshold, and $\mathrm{nonzero}(\cdot)$ is an operation employed to calculate the number of nonzero vectors in $\mathcal{B}$. In light of this, each skill corresponds to a skill-wise code $h$.
@@
0:  Initialized adaptive language semantic bank $\mathcal{B}$ with $N^{b}$ zero vectors; Initialized dynamic skill-specific latent space $\mathbf{S}=\emptyset$; Initialized low-rank space $\mathbf{W}=\emptyset$; The hyper-parameter $\delta$;
0:  language embedding $\mathbf{l}_{s}$;
0:  $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
1:  Compute cosine similarity matrix $\mathcal{C}$ between $\mathcal{B}$ and $\mathbf{l}_{s}$;
2:  if $\mathcal{C}_{max}>\delta$ then
3:     $h\xleftarrow{}\mathrm{argmax}(\mathcal{C});$
4:     Update $\mathcal{B}$ by Eq. (8);
5:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
6:  else
7:     $h\xleftarrow{}\mathrm{nonzero}(\mathcal{B}+1);$
8:     Expand $\mathcal{B}$ by Eq. (8);
9:     Randomly initialize $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
10:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
11:  end if
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times D}$ are linear projection layers, and $\mathrm{Cat}(\cdot)$ denotes the concatenation operation. $\rho$ indices the softmax function and $d$ is a scaling factor. $\mathbf{p}\in\mathbb{R}^{N^{p}\times D}$ is obtained by concatenating the current state of agent and the input voxelized observation. $\mathbf{e}\in\mathbb{R}^{N^{e}\times D}$ is the encoded embeddings of the input language instruction. To this end, relying on the skill-wise code $h$, NBAgent can continually encode the novel multi-modal input from a skill-specific latent space, which is beneficial to learn skill-specific knowledge from latent space.
@@
Figure 2: Visualization of prediction results on various manipulation skills between Ours, Ours-w/oSRD, Ours-w/oSEP&SRD, ER and Fine-Tuning.
@@
$\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{SSR}}+\lambda_{2}\mathcal{L}_{\mathrm{SRD}}.$ (11)
4 Experiments
4.1 Implementation Details
Dataset. By following PerAct [39], we conduct our experiments on RLBench [22] and simulate robot learning in CoppelaSim [34]. To simulate the working scenarios of NBRL, we design two NBRL benchmark datasets, called Kitchen and Living Room. Specifically, Kitchen is constructed by gathering 10 manipulation skills pertinent to kitchen environments, and Living Room consists of 12 manipulation skills associated with living room scenarios. Each manipulation skill includes a training set of 20 episodes and a test set of 25 episodes. Furthermore, these skills involve various variations encompassing randomly sampled attributes such as colors, sizes, counts, placements, and object categories, resulting in a total of 101 distinct variations. In this paper, we use fixed skill ID 1-10 and 1-12 to indicate each skill in different scenarios. The correspondence between each skill ID and the details of skills are provided in Appendix A.1.
Baselines. We conduct the comprehensive evaluation between our NBAgent and the following four methods: PerAct [39], GNFactor [53], Fine-Tuning and ER [7]. PerAct and GNFactor joint train all manipulation skills within one-stage training, for two datasets respectively, referred to as the upperbound. Fine-Tuning achieves continual learning by fine-tuning all parameters in perceiver model on novel skills. ER randomly stores old skills data to memory buff and replay them when detecting novel skills.
Table 1: Comparisons of success rate (%) on Kitchen and Living Room. Red and Blue represents the highest results and runner-up. The number of steps equals the sum of quantity of base skill learning task and continual skill learning tasks.
@@ 4.1 Implementation Details
[…]
@@
ER 56.0 25.6 40.8 50.0 3.2 53.6 29.6 41.6 49.7 9.8 43.8 31.1 37.4 42.2 7.7 40.7 30.2 35.4 38.1 17.6
Ours-w/oSEP&SRD 56.0 26.7 41.3 49.1 0.8 56.8 30.4 43.6 50.2 7.1 42.0 33.1 37.6 45.3 12.6 38.4 35.6 37.0 40.9 16.5
Ours-w/oSRD 41.1 38.1 39.6 49.8 18.9 48.0 41.6 44.8 53.9 14.7 41.1 34.4 37.8 45.9 15.6 40.4 39.1 39.8 43.3 19.1
@@
Training Details. In NBRL, following continual learning setting [12], we assume that the agent undergoes initial learning through a set of manipulation skills, referred to as the base skill learning task. Subsequently, skills are added incrementally in steps, characterized as continual skill learning tasks. To valuate the robustness of NBAgent in addressing NBRL problem, we set up four combinations of base skills and new skills: 5-5, 5-1, 6-3, and 6-2, which are highlighted in bold in Tab. 1. For instance, ’5-1’ indicates that the agent is initially trained on a base skill learning task comprising the first five manipulation skills in the skill ID. Subsequently, the agent undergoes five continual skill learning tasks, each involving a new skill, ultimately learning a total of 10 skills in the Kitchen dataset. In addition, the LAMB [51] optimizer is applied for all methods with a initial learning rate of $5.0\times 10^{-4}$ and a batch size of 2. We utilize 100K training iterations for PerAct and GNFactor, 80K for base task training, 20K for incremental task training. We randomly store a fixed 4 episodes of each old skill in $\mathcal{M}$ for ER and Ours. Additional training details are available in Appendix A.2.
@@
4.2 Comparison Performance

### v3

@@ III-C Skill-Shared Representation Distillation Module
where $\hat{B}=\sum_{j=1}^{B}\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}}$, $\rho$ indices the softmax function. $\tau$ is a hyper-parameter represents distillation temperature.
III-D Skill-Specific Evolving Planner
@@
Figure 3: Prediction examples on RLBench [24]. For qualitative evaluation, we visualize three key frames from each manipulation task across all methods.
@@
where $\mathcal{B}\in\mathbb{R}^{N^{b}\times D^{s}}$ is initialized by $N^{b}$ zero vectors. $\mathcal{C}\in\mathbb{R}^{N^{b}}$ is the cosine similarity matrix between the sentence information $\mathbf{l}_{s}$ and each vector in $\mathcal{B}$. if $\mathcal{C}_{max}>\delta$, we set $h=\mathrm{argmax}(\mathcal{C})$; otherwise, we set $h=\mathrm{nonzero}(\mathcal{B})+1$ and $\mathcal{C}_{max}=1$, where $\delta=0.8$, and $\mathrm{nonzero}(\cdot)$ is an operation employed to calculate the number of nonzero vectors in $\mathcal{B}$. In light of this, each skill corresponds to a skill-wise code $h$.
Initialize: Adaptive language semantic bank $\mathcal{B}$ with $N^{b}$ zero vectors; Dynamic skill-specific latent space $\mathbf{S}=\emptyset$; Low-rank space $\mathbf{W}_{r}=\emptyset$; Hyper-parameter $\delta$;
$\triangleright$ Input: language embedding $\mathbf{l}_{s}$;
Compute cosine similarity matrix $\mathcal{C}$ between $\mathcal{B}$ and $\mathbf{l}_{s}$;
if $\mathcal{C}_{max}>\delta$ then
 $h\xleftarrow{}\mathrm{argmax}(\mathcal{C});$
  Update $\mathcal{B}$ by Eq. (11);
 Return: $\mathbf{S}[h,:],\mathbf{W}_{r}[h,:]$;
else
 $h\xleftarrow{}\mathrm{nonzero}(\mathcal{B})+1;$
   Expand $\mathcal{B}$ by Eq. (11);
   Randomly initialize $\mathbf{S}[h,:],\mathbf{W}_{r}[h,:]$;
 Return: $\mathbf{S}[h,:],\mathbf{W}_{r}[h,:]$;
$\triangleright$ Output: $\mathbf{S}[h,:],\mathbf{W}_{r}[h,:]$.
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times D}$ are linear projection layers, and $\mathrm{Cat}(\cdot)$ denotes the concatenation operation. $d$ is a scaling factor. $\mathbf{p}\in\mathbb{R}^{N^{p}\times D}$ is obtained by concatenating the current state of agent and the input voxelized observation. $\mathbf{e}\in\mathbb{R}^{N^{e}\times D}$ is the encoded embeddings of the input language instruction. To this end, relying on the skill-wise code $h$, NBAgent can continually encode the novel multi-modal input from a skill-specific latent space, which is beneficial to learn skill-specific knowledge from latent space.
Figure 4: Visualization of manipulation skills in dataset Kitchen, consisting of 10 manipulation skills pertinent to kitchen environments.
Figure 5: Visualization of manipulation skills in dataset Living Room, including 12 manipulation skills associated with living room scenarios.
@@
$\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{SSR}}+\lambda_{2}\mathcal{L}_{\mathrm{SRD}},$ (14)
where $\lambda_{2}$ represents hyper-parameter. $\Theta^{m}_{p},\Theta^{m}_{n}$ are optimized via $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{SSR}}$ for the first task, and trained via $\mathcal{L}_{\mathrm{Total}}$ in Eq. (14) when $m\geq 2$.
TABLE I: Skill Details in Kitchen & Living Room
ID Skill in Kitchen Language Example VA. Avg. Kf
1 close microwave “close microwave” 1 2.3
2 meat off grill “take the steak off the grill” 2 6.0
3 open grill “open the grill” 1 4.5
4 open wine bottle “open wine bottle” 1 3.3
5 pick up cup “pick up the red cup” 20 3.8
6 turn tap “turn left tap” 2 3.0
7 place wine “stack the wine bottle” 3 6.1
8 put knife on “put the knife on the board” 1 5.1
9 stack wine “stack wine bottle” 1 7.4
10 take plate off “take plate off colored rack” 3 6.3
ID Skill in Living Room Language Example VA. Avg. Kf
1 close door “close the door” 1 4.8
2 close laptop lid “close laptop lid” 1 6.0
3 hang frame “hang frame on hanger” 1 5.3
4 lamp on “turn on the light” 1 3.5
5 open drawer “open the bottom drawer” 3 4.8
6 open window “open left window” 1 6.0
7 push buttons “push the white button” 18 4.8
8 put item in drawer “put the item in the top drawer” 3 14.3
9 put rubbish in bin “put rubbish in bin” 1 5.0
10 sweep to dustpan “sweep dirt to the tall dustpan” 2 6.2
11 take usb out “take usb out of computer” 1 3.3
12 water plants “pour some water on the plant” 1 6.3
IV Experiments
We conduct experiments to investigate how NBAgent leverages both skill-shared and skill-specific knowledge learning to address the NBRL problem on two simulation datasets. To this end, we design five experimental paradigms: (1) comparison performance on RLBench using simulation data to validate overall learning capability (Sec. IV-B and Sec. IV-C), (2) ablation studies on key modules and hyperparameters to quantify their individual contributions (Sec. IV-D), (3) task complexity analysis to examine performance under varying skill complexities (Sec. IV-E), and (4) computational complexity analysis to assess training and inference efficiency (Sec. IV-F).
IV-A Implementation Details
Dataset. By following PerAct [53], we conduct our experiments on RLBench [24] and simulate robot learning in CoppelaSim [47]. Here, language annotations are obtained through manually designed templates, while key frames are extracted by monitoring the robot’s state and task-related events during simulation. To simulate the working scenarios of NBRL, we design two NBRL benchmark datasets, called Kitchen and Living Room. Specifically, as shown in Figs. 4-5 and Tab. I, Kitchen dataset is constructed by gathering 10 manipulation skills pertinent to kitchen environments, and Living Room dataset consists of 12 manipulation skills associated with living room scenarios. Each manipulation skill includes a training set of 20 episodes and a test set of 25 episodes. Furthermore, these skills involve various variations encompassing randomly sampled attributes such as colors, sizes, counts, placements, and object categories, resulting in a total of 69 distinct variations. In this section, we employ a fixed skill identification scheme, where skills are uniquely indexed within the ranges of 1-10 and 1-12 for distinct operational scenarios. A comprehensive summary of these skills, including their functional descriptions, associated key frame statistics and variation quality, is presented in Tab. I. Specifically, the ID column denotes the unique identifier for each skill, while the Avg. Kf metric quantifies the mean number of key frames observed across 20 training episodes. Additionally, Va. denoting the number of variations per skill, provides a summary of the dataset and reflects the diversity and complexity of the contained scenarios.
TABLE II: Comparisons of success rate (%) on Kitchen and Living Room. Red and Blue represents the highest results and runner-up. The number of steps equals the sum of quantity of base skill learning task and continual skill learning tasks.
@@ IV-A Implementation Details
PerAct [53] (CoRL’2023) 58.9 30.1 44.5 $\mathrm{-}$ $\mathrm{-}$ 58.9 30.1 44.5 $\mathrm{-}$ $\mathrm{-}$ 34.7 27.3 31.0 $\mathrm{-}$ $\mathrm{-}$ 34.7 27.3 31.0 $\mathrm{-}$ $\mathrm{-}$
GNFactor [70] (CoRL’2023) 56.3 32.3 44.3 $\mathrm{-}$ $\mathrm{-}$ 56.3 32.3 44.3 $\mathrm{-}$ $\mathrm{-}$ 48.0 32.0 40.0 $\mathrm{-}$ $\mathrm{-}$ 48.0 32.0 40.0 $\mathrm{-}$ $\mathrm{-}$
@@
ER [8] (ECCV’2018) 56.0 25.6 40.8 50.0 3.2 53.6 29.6 41.6 49.7 9.8 43.8 31.1 37.4 42.2 7.7 40.7 30.2 35.4 38.1 17.6
LWF [35] (TPAMI’2017) 57.6 30.4 44.0 51.0 1.8 52.0 30.2 41.1 50.4 11.5 40.0 32.7 36.3 42.9 13.5 43.2 31.4 37.2 41.0 15.8
iCaRL [46] (CVPR’2017) 42.4 34.4 38.4 49.8 15.2 54.0 33.4 43.7 51.0 9.8 42.0 32.0 37.0 43.1 11.8 44.0 30.8 37.4 41.5 15.4
PODNet [14] (ECCV’2020) 52.8 30.4 41.6 50.6 10.8 53.6 31.2 42.4 53.2 9.6 43.4 32.7 38.1 43.9 11.6 44.7 32.0 38.4 42.2 12.7
C-LoRA [54] (TMLR’2024) 48.0 32.8 40.4 40.2 13.6 49.6 35.2 42.4 52.5 10.3 42.0 39.3 40.7 44.6 9.6 40.7 34.7 37.7 41.8 14.7
M2Distill [49] (ICRA’2025) 55.5 30.9 43.2 51.0 8.3 53.1 34.4 43.7 52.8 10.7 45.8 36.5 41.1 45.8 10.2 43.6 28.0 38.4 40.8 15.1
@@
Figure 6: Comparisons of forgetting rate (i.e., For.) (%) under the settings of 5+1 (6 steps) and 5+1 (12 steps).
Baselines. We conduct the comprehensive evaluation between our NBAgent and the following two multi-task behavior-cloning methods: PerAct [53], GNFactor [70], and five continual learning methods in behavior-cloning agent setting: Fine-Tuning, ER [9], LWF [35], iCaRL [46], PODNet [14] and C-LoRA [54]. More specifically, both PerAct [53] and GNFactor[70] joint train all manipulation skills within one-stage training, for two datasets respectively, referred to as the upperbound. Fine-Tuning enables continual learning by updating all parameters of the Perceiver model for novel skills. ER [8] randomly stores old skills to memory buff and replay them when learning novel skills. LWF [35] applies knowledge distillation between current and old models using replayed samples. iCaRL [46] introduces memory management to select and update stored episodes in the memory buffer. C-LoRA [54] employs an orthogonal regularization loss to prevent catastrophic forgetting during task-specific LoRA adaptation. M2distill [49] preserves a consistent latent space across vision, language, and action modalities to mitigate catastrophic forgetting.
TABLE III: Comparison results on Kitchen under the setting of 5-5.
Skill ID 1 2 3 4 5 6 7 8 9 10 1-5 6-10 All. Imp.
PerAct [53] (CoRL’2023) 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 58.9 30.1 44.5 $\Uparrow$ 0.4
GNFactor [70] (CoRL’2023) 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 56.3 32.3 44.3 $\Uparrow$ 0.6
Fine-Tuning 10.7 0.0 33.3 33.3 1.3 49.3 6.7 14.7 5.3 56.0 15.7 26.4 21.1 $\Uparrow$ 23.8
ER [8] (ECCV’2018) 96.0 72.0 56.0 18.7 37.3 45.3 0.0 14.7 1.3 66.7 56.0 25.6 40.8 $\Uparrow$ 4.1
LWF [35] (TPAMI’2017) 97.3 72.0 54.7 33.3 33.3 48.0 6.7 34.7 4.0 56.0 57.6 30.4 42.8 $\Uparrow$ 2.1
iCaRL [46] (CVPR’2017) 60.0 41.3 77.3 22.7 32.0 61.3 2.7 21.3 10.7 81.3 42.4 34.4 38.4 $\Uparrow$ 6.5
PODNet [14] (ECCV’2020) 93.3 74.7 42.7 24.0 28.0 41.3 9.3 32.0 6.7 64.0 52.8 30.4 41.6 $\Uparrow$ 3.3
C-LoRA [54] (TMLR’2024) 74.7 40.0 56.0 33.3 36.0 49.3 17.3 24.0 12.0 61.3 48.0 32.8 40.4 $\Uparrow$ 4.5
M2Distill [49] (ICRA’2025) 84.0 70.7 56.0 29.3 37.3 52.0 0.0 29.3 9.3 64.0 55.5 30.9 43.2 $\Uparrow$ 1.7
Ours 94.7 52.0 38.7 44.0 38.7 64.0 2.7 34.7 9.3 70.7 53.6 36.3 44.9 $\mathrm{-}$
TABLE IV: Comparison results on Kitchen under the setting of 5-1.
Skill ID 1 2 3 4 5 6 7 8 9 10 1-5 6-10 All Imp.
PerAct [53] (CoRL’2023) 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 58.9 30.1 44.5 $\Uparrow$ 1.5
GNFactor [70] (CoRL’2023) 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 56.3 32.3 44.3 $\Uparrow$ 1.7
Fine-Tuning 17.3 0.0 0.0 0.0 0.0 8.0 0.0 0.0 9.3 64.0 3.2 20.0 9.7 $\Uparrow$ 36.3
ER [8] (ECCV’2018) 90.7 54.7 37.3 54.7 28.0 60.0 5.3 0.0 24.0 62.7 53.6 29.6 41.6 $\Uparrow$ 4.4
LWF [35] (TPAMI’2017) 88.0 50.7 42.7 41.3 37.3 53.3 10.7 14.7 9.3 64.0 52.0 30.2 41.1 $\Uparrow$ 4.9
iCaRL [46] (CVPR’2017) 89.3 54.7 56.0 32.0 38.7 49.3 8.0 29.3 14.7 68.0 54.0 33.4 43.7 $\Uparrow$ 2.3
PODNet [14] (ECCV’2020) 92.0 57.3 53.3 32.0 34.7 52.0 14.7 24.0 13.3 50.7 53.6 31.2 42.4 $\Uparrow$ 3.6
C-LoRA [54] (TMLR’2024) 80.0 49.3 50.7 29.3 40.0 69.3 18.7 16.0 14.7 57.3 49.6 35.2 42.4 $\Uparrow$ 3.6
M2Distill [49] (ICRA’2025) 88.0 56.0 49.3 32.0 40.0 70.7 14.7 17.3 14.7 54.7 53.1 34.4 43.7 $\Uparrow$ 2.3
Ours 92.0 61.3 44.0 34.7 41.3 76.0 17.3 26.7 14.7 52.0 54.4 37.6 46.0 $\mathrm{-}$
@@
Training Details. In NBRL, following continual learning setting [14], we assume that the agent undergoes initial learning through a set of manipulation skills, referred to as the base skill learning task. Subsequently, skills are added incrementally in steps, characterized as continual skill learning tasks. To valuate the robustness of NBAgent in addressing NBRL problem, we set up four combinations of base skills and new skills: 5-5, 5-1, 6-3, and 6-2, which are highlighted in bold in Tab. II. For instance, “5-1” indicates that the agent is initially trained on a base skill learning task comprising the first five manipulation skills in the skill ID. Subsequently, the agent undergoes five continual skill learning tasks, each involving a new skill, ultimately learning a total of 10 skills in the Kitchen dataset. In NBRL-R dataset, we adopt an NBRL setting denoted as “4-1”, where the agent first learns the initial four skills listed in Tab LABEL:tab:_simpler_details and then continually learns one new skill over 3 steps. Additionally, we use RT-1 [5] as the base behavior-cloning agent and apply the proposed NBAgent, along with other comparison methods, to perform NBRL. To evaluate the long-term learning capability of NBAgent against other methods, we design a novel long-term learning setting, denoted as 5-1 (12 steps). In this setting, the agent is initially trained on the first five manipulation skills from the Kitchen dataset and subsequently undergoes 11 continual learning tasks, each introducing a new skill drawn from the remaining five Kitchen skills and the first six Living Room skills.
In addition, the LAMB [67] optimizer is applied for all methods with a initial learning rate of $5.0\times 10^{-4}$ and a batch size of 2. We utilize 100K training iterations for PerAct and GNFactor, 80K for base task training, 20K for incremental task training. We store a fixed 4 episodes of each old skill in $\mathcal{M}$ for all methods. We configure the hyper-parameters as $\alpha=0.1$ in Eq. (III-B), $\delta=0.8$ in SEP module, with $\lambda_{1}=0.1$ and $\lambda_{2}=0.2$ to balance $\mathcal{L}_{S}$ in Eq. (7) and $\mathcal{L}_{SRD}$ in Eq. (14).
@@
TABLE V: Comparison results on Living Room under the setting of 6-3.
Skill ID 1 2 3 4 5 6 7 8 9 10 11 12 1-6 7-12 All Imp.
PerAct [53] (CoRL’2023) 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 34.7 27.3 31.0 $\Uparrow$ 12.6
GNFactor [70] (CoRL’2023) 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 48.0 32.0 40.0 $\Uparrow$ 3.6
Fine-Tuning 8.0 18.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 68.0 100.0 10.7 4.4 29.8 17.1 $\Uparrow$ 25.6
ER [8] (ECCV’2018) 10.7 58.7 24.0 81.3 76.0 12.0 5.3 4.0 78.7 8.0 89.3 1.3 43.8 31.1 37.4 $\Uparrow$ 6.2
LWF [35] (TPAMI’2017) 16.0 62.7 9.3 82.7 52.0 17.3 4.0 9.3 52.0 33.3 84.0 14.7 40.0 32.7 36.3 $\Uparrow$ 7.3
iCaRL [46] (CVPR’2017) 29.3 80.0 21.3 74.7 36.0 10.7 4.0 2.7 68.0 1.3 96.0 20.0 42.0 32.0 37.0 $\Uparrow$ 6.6
PODNet [14] (ECCV’2020) 13.3 56.0 26.7 78.7 61.3 25.3 1.3 6.7 72.0 8.0 82.7 22.7 43.4 32.7 38.1 $\Uparrow$ 5.5
C-LoRA [54] (TMLR’2024) 16.0 53.3 25.3 74.7 52.0 32.0 6.7 25.3 72.0 14.7 82.7 33.3 42.0 39.3 40.7 $\Uparrow$ 2.9
M2Distill [49] (ICRA’2025) 13.3 74.7 28.0 80.0 54.7 24.0 10.7 13.3 62.7 14.7 88.0 29.3 45.8 36.5 41.1 $\Uparrow$ 2.5
Ours 17.3 74.7 9.3 82.7 62.7 22.7 22.7 9.3 85.3 9.3 98.7 28.0 44.9 42.2 43.6 $\mathrm{-}$
TABLE VI: Comparison results on Living Room under the setting of 6-2.
Skill ID 1 2 3 4 5 6 7 8 9 10 11 12 1-6 7-12 All Imp.
PerAct [53] (CoRL’2023) 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 34.7 27.3 31.0 $\Uparrow$ 9.6
GNFactor [70] (CoRL’2023) 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 48.0 32.0 40.0 $\Uparrow$ 0.6
Fine-Tuning 0.0 37.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 96.0 18.7 6.2 19.1 12.7 $\Uparrow$ 27.9
ER [8] (ECCV’2018) 4.0 89.3 17.3 80.0 38.7 14.7 20.0 2.7 60.0 4.0 90.7 4.0 40.7 30.2 35.4 $\Uparrow$ 5.2
LWF [35] (TPAMI’2017) 12.0 53.3 45.3 60.0 61.3 33.3 4.0 10.7 73.3 5.3 84.0 5.3 43.2 31.4 37.2 $\Uparrow$ 3.4
iCaRL [46] (CVPR’2017) 17.3 52.0 45.3 68.0 64.0 17.3 2.7 24.0 70.7 10.7 72.0 4.0 44.0 30.8 37.4 $\Uparrow$ 3.2
PODNet [14] (ECCV’2020) 10.7 54.7 52.0 57.3 60.0 32.0 9.3 14.7 70.7 14.7 72.0 13.3 44.7 32.0 38.4 $\Uparrow$ 2.2
C-LoRA [54] (TMLR’2024) 16.0 48.0 45.3 62.7 52.0 18.7 17.3 33.3 70.7 14.7 62.7 10.7 40.7 34.7 37.7 $\Uparrow$ 2.9
M2Distill [49] (ICRA’2025) 16.0 54.7 48.0 81.3 61.3 32.0 9.3 9.3 64.0 10.7 70.7 4.0 43.6 28.0 38.4 $\Uparrow$ 2.2
Ours 8.0 61.3 32.0 84.0 58.7 30.7 4.0 17.3 80.0 6.7 98.7 5.3 45.8 35.3 40.6 $\mathrm{-}$
TABLE VII: Ablation studies on the Kitchen dataset under the settings of 5-5 and 5-1.
Variants Base Base+SSR Base+SSR Base+SSR   Ours
  +SEP   +SRD
Modules SSR
SEP
SRD
5-5 $(2~\mathrm{steps})$ 1-5 48.6 56.0 41.1 54.5 53.6
6-10 26.4 26.7 38.1 28.6 36.3
ALL 37.5 41.3 39.6 41.6 44.9
Avg. 45.8 49.1 49.8 51.2 52.5
For. 18.4 0.8 18.9 5.8 6.4
5-1 $(6~\mathrm{steps})$ 1-5 52.0 56.8 48.0 57.2 54.4
6-10 27.8 30.4 41.6 30.8 37.6
ALL 40.3 43.6 44.8 44.0 46.0
Avg. 49.4 50.2 53.9 53.1 55.2
For. 10.6 7.1 14.7 6.5 8.9
TABLE VIII: Ablation studies on the Living Room dataset under the settings of 6-3 and 6-2.
Variants Base Base+SSR Base+SSR Base+SSR   Ours
  +SEP   +SRD
Modules SSR
SEP
SRD
6-3 $(3~\mathrm{steps})$ 1-6 41.8 42.0 41.1 46.6 44.9
7-12 31.6 33.1 34.4 32.8 42.2
ALL 36.7 37.6 37.8 39.7 43.6
Avg. 42.4 45.3 45.9 46.4 47.6
For. 11.9 12.6 15.6 6.8 7.7
6-2 $(4~\mathrm{steps})$ 1-6 40.3 38.4 40.4 44.7 45.8
7-12 30.6 35.6 39.1 33.8 35.3
ALL 35.5 37.0 39.8 39.2 40.6
Avg. 39.2 40.9 43.3 42.0 43.5
For. 14.1 16.5 19.1 11.6 10.9
IV-B Comparison Performance on RLbench